OpenStudy (anonymous):

What is the standard form of the equation of the conic given by 2x^2 + 2y^2 + 4x - 12y - 22 = 0 ?

OpenStudy (anonymous):

\[2x ^{2} + 2y ^{2} + 4x - 12y - 22 = 0\]

OpenStudy (anonymous):

The choices are: a. \[\frac{ (x+1)^{2} }{ 21 } - \frac{ (y-3)^{2} }{ 21 } = 1\] b. \[\frac{ (x+1)^{2} }{ 21 } + \frac{ (y-3)^{2} }{ 21 } = 1\] c. \[\frac{ (x-3)^{2} }{ 21 } + \frac{ (y+1)^{2} }{ 21 } = 1\] d. \[\frac{ (x-1)^{2} }{ 7 } + \frac{ (y+3)^{2} }{ 3 } = 1\]

OpenStudy (anonymous):

Did you ever get the answer?